OpenStudy (anonymous):

partials fractions help please... s^2(s+2)/(s^2-16)

hartnn (hartnn):

let me guess, for inverse laplace transform ?

hartnn (hartnn):

factor s^2-16 first

OpenStudy (anonymous):

lol yeah

hartnn (hartnn):

a^2-b^2 = (a+b)(a-b)

OpenStudy (anonymous):

so s^3(s+2)/(s^2+4)(s^2-4)

hartnn (hartnn):

is it s^4 -16 ? O.o

OpenStudy (anonymous):

yeah sorry

hartnn (hartnn):

ok, no problem :)

hartnn (hartnn):

for denominator of \(s^2 \pm a^2\) or any other quadratic, we use \(\Large \dfrac{As+B}{quadratic}\)

hartnn (hartnn):

so, \(\dfrac{s^2(s+2)}{()()}= \dfrac{As+B}{()} +\dfrac{Cx+D}{()}\) i feel lazy to fill denominators hope you can fill it :P

OpenStudy (anonymous):

yeah so far so good followoing along on paper

OpenStudy (anonymous):

(as+d)(s^2-4) + (bx+c)(s^2+4) = s^3 = 2s^2

OpenStudy (anonymous):

I mean = s^3 + 2s^2

hartnn (hartnn):

expand left side so that you can compare the co-efficients

OpenStudy (anonymous):

as^3 + bs^3 = 1 and as^2 + cs^2 =2

hartnn (hartnn):

you mean as^3+bs^3 = s^3 giving a+b = 1 ?

OpenStudy (anonymous):

yes I have a+b = 1

hartnn (hartnn):

start with easier terms -4d+4c = 0 d=c

OpenStudy (anonymous):

and a = b from -4a +4b = 0

OpenStudy (anonymous):

got that

OpenStudy (anonymous):

 @hartnn

hartnn (hartnn):

but the question has s^4 -16 read comments please

OpenStudy (anonymous):

@Hanishala I typed in wrong term to start with );

OpenStudy (anonymous):

ok so I also have ds^2 + cs^2 = 2s^2 correct??

OpenStudy (anonymous):

sorry i didnot see it

OpenStudy (anonymous):

its ok

hartnn (hartnn):

(as+d)(s^2-4) + (bs+c)(s^2+4) = s^3 + 2s^2 d+c =2 yes correct o on

hartnn (hartnn):

*go on

OpenStudy (anonymous):

so we have d - 1, c =1, a =1/2, b = 1/2 right?

OpenStudy (anonymous):

d= 1, c = 1, a =1/2 b = 1/2

OpenStudy (anonymous):

correct?

hartnn (hartnn):

let me cross check again

hartnn (hartnn):

we made an error from the beginning :(

OpenStudy (anonymous):

where at?

hartnn (hartnn):

sorry s^2-4 can be factored further

OpenStudy (anonymous):

ok

hartnn (hartnn):

which means s+2 gets cancelled!

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

you can make s^2 -4 = (s-2) +(s+2 ) and continue

hartnn (hartnn):

s^3/ (s^2+4)(s-2) = A/(s-2) + (Bs+C)/(s^2+4)

hartnn (hartnn):

s^3 = A(s^2+4) + (Bs+C)(s-2) this is even easier! :)

hartnn (hartnn):

to find A, plug in s=2 :)

OpenStudy (anonymous):

wouldnt it be s^2 on left side now?

hartnn (hartnn):

yes, sorry typo

hartnn (hartnn):

but i think you can find A,B,C try ?

OpenStudy (anonymous):

Yeah let me try (;

OpenStudy (anonymous):

Thank you both.

hartnn (hartnn):

so what values did u get ?